# NeRAF: 3D Scene Infused Neural Radiance and Acoustic Fields

### v3

@@ Abstract
Sound plays a major role in human perception. Along with vision, it provides essential information for understanding our surroundings. Despite advances in neural implicit representations, learning acoustics that align with visual scenes remains a challenge. We propose NeRAF, a method that jointly learns acoustic and radiance fields. NeRAF synthesizes both novel views and spatialized room impulse responses (RIR) at new positions by conditioning the acoustic field on 3D scene geometric and appearance priors from the radiance field. The generated RIR can be applied to auralize any audio signal. Each modality can be rendered independently and at spatially distinct positions, offering greater versatility. We demonstrate that NeRAF generates high-quality audio on SoundSpaces and RAF datasets, achieving significant performance improvements over prior methods while being more data-efficient. Additionally, NeRAF enhances novel view synthesis of complex scenes trained with sparse data through cross-modal learning. NeRAF is designed as a Nerfstudio module, providing convenient access to realistic audio-visual generation. Project page: https://amandinebtto.github.io/NeRAF
@@ 2 Related works
RIR have various applications, such as enabling immersive and spatialized audio for AR and VR, performing de-reverberation or providing insights about room acoustics. However, RIR collection in real-world environment is time consuming and needs specialized hardware. Consequently, synthesizing them has been a longstanding research topic. Traditional methods rely on simulated approaches such as wave-based (Gumerov & Duraiswami, 2009; Thompson, 2006; Raghuvanshi et al., 2009; Bilbao et al., 2019) or geometric methods (Savioja & Svensson, 2015; Schissler & Manocha, 2016; Krokstad et al., 1968; Vorländer, 1989; Cao et al., 2016; Sprunck et al., 2022). While these methods effectively simulate sound propagation, wave-based methods face computation complexity and geometric methods struggle to simulate low-frequency acoustic phenomena such as interference and diffraction. Recent works have proposed to leverage the modeling ability of neural networks to learn room acoustics and estimate RIRs (Majumder et al., 2022; Luo et al., 2022; Su et al., 2022; Ratnarajah et al., 2021; 2022).
@@ 3 Method
NeRAF generates both RGB images and binaural RIRs for any camera and microphone-source position and orientation. It achieves this by jointly learning neural radiance and acoustic fields. Although trained together, the audio and visual components can be used independently.
The NeRAF pipeline, shown in Figure 2, consists of three modules: NeRF, the grid sampler, and NAcF. The NeRF learns the radiance field and generates RGB images. The grid sampler queries NeRF’s radiance field with each voxel’s center, filling the 3D grid with color and density. NAcF then renders binaural RIRs, conditioned by the implicit appearance and geometric 3D priors derived from the features extracted from this 3D grid.
@@ 3.2 Neural Radiance Field
NeRF shoots rays $\textbf{r}(t)=\textbf{o}+t\textbf{d}$ through image pixels from camera origin o between the near $t_{near}$ and far $t_{far}$ boundaries of the scene. 3D points $\textbf{X}_{n}=\textbf{o}+t_{n}\textbf{d}$ are sampled along the ray and, along with d used as input of two MLPs. The first one maps $\textbf{X}_{n}$ to the view-independent density $\sigma_{n}$ and a corresponding feature vector. The second MLP takes the feature vector and d to produce view-dependent color $\textbf{c}_{n}$. The final color $C(\textbf{r})$ of an image pixel are rendered using volume rendering equations:
[…]
@@ 3.3 Grid Sampler
For color, we apply the volume rendering equation to a ray containing a single point, $\textbf{X}_{vi}$, with each viewing direction and average the values across these directions to obtain Ĉ$_{vi}$. For density, we compute the alpha compositing value $\alpha=1-\exp(-\sigma_{vi}\delta)$ where $\delta$ is a chosen small value. We populate the grid with these values and the 3D coordinates of the voxels, resulting in a 7-channel representation.
@@ 3.4 Neural Acoustic Field
The goal of our neural acoustic field (NAcF), presented in Figure 4, is to learn a continuous neural representation of the scene’s acoustics. To this end, NAcF maps microphone coordinates $\textbf{X}_{m}=(x_{m},y_{m},z_{m})$ and orientations $\textbf{d}_{m}=(\theta_{m},\phi_{m})$ along with source position $\textbf{X}_{s}=(x_{s},y_{s},z_{s})$ and orientation $\textbf{d}_{s}=(\theta_{s},\phi_{s})$ to the short-time Fourier transform (STFT) of a room impulse response:
@@
We learn RIR through their STFT representation, resulting in $\text{STFT}\in\mathbb{R}^{C\times F\times T}$, where C, F and T denote the number of channels, frequency bins, and time bins, respectively. This is motivated by the fact that the time-frequency representation is smoother and more suitable to neural network prediction than the time domain (Luo et al., 2022). Thus, NAcF learns to infer log-magnitude STFTs. To better capture perceptual effects such as reverberation, we query NAcF with STFT time bins $t$. Consequently, NAcF outputs the $F$ frequency bins corresponding to the time query. We normalize $t$ between $[0,1]$ and encode it with positional encoding. Similar to NeRF (Mildenhall et al., 2020) pose, microphone and source coordinates as well as direction are also embedded to a higher dimensional space using respectively positional encoding and spherical harmonic encoding.
@@ 3.5 Learning Objective
We do not modify the NeRF loss function $\mathcal{L}_{V}$. It usually consists in an MSE loss between ground-truth pixel color $C(\textbf{r})$ and the rendered one $\hat{C}(\textbf{r})$, and complementary losses, $\mathcal{L}_{\text{comp}}$, specific to each NeRF.
@@
where $\hat{s}$ is the predicted STFT, $s$ the ground-truth STFT, .$\lVert.\rVert_{F}$ denotes the Frobenius norm, .$\lVert.\rVert_{2}$ the L2 norm and $\epsilon=10^{-3}$. The values of our loss hyperparameters are found empirically and given in Appendix C. The spectral convergence loss emphasizes on spectral peaks helping especially in early phases of training while the spectral loss focus on small amplitude components which tends to be more important towards the later phases of training. Our loss differs from (Yamamoto et al., 2019; 2020) as we found out that the MSE spectral loss leads to a better trade-off between metrics than the L1 spectral loss (Su et al., 2022) and the MSE loss on magnitude STFT (Luo et al., 2022; Liang et al., 2023a). NeRAF results with those three losses are presented in Section 4.3. The complete learning objective can be resumed as:
@@ 4 Experiments
First, we evaluate NeRAF’s performance against related works (Luo et al., 2022; Su et al., 2022; Liang et al., 2023a; b) on SoundSpaces and RAF datasets. Then, we demonstrate that cross-modal learning of the acoustic and radiance fields is also beneficial to vision, resulting in enhanced results for large complex scenes trained with very sparse observations. Next, we conduct a few-shot experiment to assess the effectiveness of our method when trained with a limited number of audio recordings. Finally, we investigate the impact of the 3D grid and our loss combination on NeRAF’s performance. Video examples showing NeRAF audio-visual generation on real and synthetic scenes are presented in the supplementary material.
4.1 Datasets
SoundSpaces.
RAF.
RWAVS.
@@ 4.2 Evaluation
[…]
@@
For RAF, we follow previous works experimental setup: we keep $80\%$ of data for training and $20\%$ for evaluation. Audios are cut to $0.32$ s, resampled to $48$ kHz and processed into STFT using $N_{\text{FFT}}=1024$, $W=512$ and $H=256$. For the vision part, we randomly select one third of the data from cameras 20 to 25. Nerfstudio automatically keep 90% of them for training and 10% for evaluation.
To ensure statistically significant results, we conduct each experiment 6 times and average results. More details on the implementation and architecture are presented in Appendix C, Appendix B.
@@ Results.
As SoundSpaces is a simulated dataset that presents limitations in term of visual and acoustics realism, we also present results on RAF in Table 1 (Right). NeRAF outperforms previous works best results, achieving $-6.7$% T60, $-14.2$% C50, $-4.5$% EDT and $-52.5$% STFT error. The second part of the table contains a comparison of methods using an additional energy decay loss term (Majumder et al., 2022; Liang et al., 2023b). While NACF initially relies on this term, other methods, with this new term, are denoted by ”++”. NeRAF++ still achieves state-of-the-art results, in particular for C50, EDT and STFT error. We present details about this loss in Appendix C. We showcase qualitative results of audio synthesis in Figure 5 and in Appendix I.
@@
Previous works evaluated on SoundSpaces rely on a significant amount of RIR data, which is often unavailable in real-life scenarios. Therefore, we assess the performance of our model when trained with fewer data. We conduct a few-shot experiment on one small, one medium, and one large room and average the results. As shown in Figure 7, NeRAF is better than AV-NeRF on all metrics when trained with 75% of the data. It still outperforms AV-NeRF on EDT when trained with only 25% of the data. On RAF dataset – which is a real-world dataset that already contains a more realistic amount of training data – NeRAF still outperforms the previous methods when trained with 25% less data. Results on RAF are presented in Appendix E.
@@ Conclusion.
Acknowledgments and Disclosure of Funding
We would like to thank Simon de Moreau for his help and insights. This work was supported by the French Agence Nationale de la Recherche (ANR), under grant ANR22-CE94-0003 and was performed using HPC resources from GENCI–IDRIS (Grant 2024-103789).
@@ Appendix B Architectures
The grid sampler creates a 3D grid of voxels with a given resolution. Grid coordinates are comprised between $[0,1]^{3}$ fitting the scene contraction performed by NeRF. Consequently, knowing the exact dimensions of the room is not necessary. Voxels center coordinates and 18 viewing directions are queried to NeRF with no architecture modification. The grid is filled with color and density information along with voxels coordinates.
@@
NAcF consists of 2 MLP blocks. The first MLP block takes as input the multimodal fusion of grid features, encoded positions, and directions obtained through vector concatenation. It comprises 5 layers, each followed by a Leaky ReLU activation with a slope of 0.1, and outputs a 512 intermediate representation of the acoustic field. The second MLP block takes this intermediate representation as input and predicts the $F$ frequencies corresponding to the time query, aiming to learn the microphone specifics (e.g., HRTF for binaural microphones in SoundSpaces). It consists of one separate MLP head per microphone channel. The final activation layer is a tanh function scaled between $[-10,10]$ to fit the STFT log-magnitude range.
@@ Appendix C Implementation Details
$\mathcal{L}_{A}=\lambda_{SC}\mathcal{L}_{SC}+\lambda_{SL}\mathcal{L}_{SL}+\lambda_{ED}\mathcal{L}_{ED}.\$ (10)
@@
We follow Liang et al. (2023b) formula. Given a magnitude STFT $\mathbf{M}$, we first compute its energy in each time window by calculating the magnitude’s square and aggregating it along the frequency dimension:
@@
[…]
$\mathcal{L}_{\mathrm{ED}}=||\log_{10}\mathbf{M}^{\prime\prime}_{g}-\log_{10}\mathbf{M}^{\prime\prime}_{p}||_{1}\enspace.$ (13)
@@
[…]
@@
STFT error ($\text{STFT error }(\mathbf{M},\mathbf{M}_{GT})=|\log(\mathbf{M})-\log(\mathbf{M}_{GT})|$ (17)
@@ Appendix D Scene Views
Real Acoustic Fields dataset captures high-quality and dense room impulse response data paired with multi-view images using VR-NeRF ”Eyeful Tower” camera rig (Xu et al., 2023). It contains 22 cameras at different heights and pointing at different directions. We chose cameras 20 to 25 for training NeRAF’s vision part. This choice ensures that we have sufficient overlapping views. The images have a 0.7 megapixels with resolution (i.e., 684 × 1024). For more details on those data we invite readers to refer to VR-NeRF paper. To train NeRAF, as we want to limit the number of images necessaries, we randomly sub-select one third of those images. We let Nerfstudio select 90% of them for training and 10% for evaluation.
@@ Appendix I Additional Results Visualization
Previous works (Liang et al., 2023a; b) rely on images (2D) to provide scene information to the acoustic field but we argue that conditioning on 3D information better suits the properties of sound propagation. For example, features like reverberation (linked to the room volume and surfaces) depend on complete 3D volume, not just what’s visible in front of the microphone or camera. By leveraging 3D information, acoustic fields can capture the acoustics of the scene more effectively.
In this paper, we propose to use NeRF as a convenient way for providing 3D scene priors —geometry and appearance— directly to the acoustic field without needing explicit 3D annotations, such as meshes or point clouds, which are complex to obtain.
@@
Joint training improves image generation in large, complex scenes with sparse visual observations, as shown in table 2 and fig. A6. We assess the impact of separate training of each modality. In that case, a NeRF (without NAcF) is fully trained and used to populate a 3D grid. Then, we train NAcF without the NeRF component. For smaller scenes with sufficient observations, such as office 4, there is no notable difference between vision performances for joint and separate training. Audio generation performance remains equivalent for both approaches, which is expected since the acoustic field in both setups leverages a 3D grid representation of the scene. Detailed results can be found in table A5.
@@
We provide multi-scene results when evaluated on each scene in table A6. Each grid is encoded to 1,024 features using the ResNet3D.
Table A6: Multi-scene training. Comparison of NeRAF performances when trained on a single scene against multiple scenes.

### v4

@@ Abstract
Sound plays a major role in human perception. Along with vision, it provides essential information for understanding our surroundings. Despite advances in neural implicit representations, learning acoustics that align with visual scenes remains a challenge. We propose NeRAF, a method that jointly learns radiance and acoustic fields. NeRAF synthesizes both novel views and spatialized room impulse responses (RIR) at new positions by conditioning the acoustic field on 3D scene geometric and appearance priors from the radiance field. The generated RIR can be applied to auralize any audio signal. Each modality can be rendered independently and at spatially distinct positions, offering greater versatility. We demonstrate that NeRAF generates high-quality audio on SoundSpaces and RAF datasets, achieving significant performance improvements over prior methods while being more data-efficient. Additionally, NeRAF enhances novel view synthesis of complex scenes trained with sparse data through cross-modal learning. NeRAF is designed as a Nerfstudio module, providing convenient access to realistic audio-visual generation. Project page: https://amandinebtto.github.io/NeRAF
@@ 2 Related works
RIRs have various applications, such as enabling immersive and spatialized audio for AR and VR, performing de-reverberation or providing insights about room acoustics. However, RIRs collection in real-world environment is time consuming and needs specialized hardware. Consequently, synthesizing them has been a longstanding research topic. Traditional methods rely on simulated approaches such as wave-based (Gumerov & Duraiswami, 2009; Thompson, 2006; Raghuvanshi et al., 2009; Bilbao et al., 2019) or geometric methods (Savioja & Svensson, 2015; Schissler & Manocha, 2016; Krokstad et al., 1968; Vorländer, 1989; Cao et al., 2016; Sprunck et al., 2022). While these methods effectively simulate sound propagation, wave-based methods face computation complexity and geometric methods struggle to simulate low-frequency acoustic phenomena such as interference and diffraction. Recent works have proposed to leverage the modeling ability of neural networks to learn room acoustics and estimate RIRs (Majumder et al., 2022; Luo et al., 2022; Su et al., 2022; Ratnarajah et al., 2021; 2022).
@@ 3 Method
NeRAF generates both RGB images and binaural RIRs for any camera and microphone-source position and orientation. It achieves this by jointly learning neural radiance and acoustic fields. Although trained together, the audio and visual components can be used independently. The NeRAF pipeline, shown in Figure 2, consists of three modules: NeRF, the grid sampler, and NAcF. The NeRF learns the radiance field and generates RGB images. The grid sampler queries NeRF’s radiance field with each voxel’s center, filling the 3D grid with color and density. NAcF then renders binaural RIRs, conditioned by the implicit appearance and geometric 3D priors derived from the features extracted from this 3D grid.
@@ 3.2 Neural Radiance Field
NeRF shoots rays $\textbf{r}(t)=\textbf{o}+t\textbf{d}$ through image pixels from camera origin o between the near $t_{\text{near}}$ and far $t_{\text{far}}$ boundaries of the scene. 3D points $\textbf{X}_{n}=\textbf{o}+t_{n}\textbf{d}$ are sampled along the ray and, along with d used as input of two MLPs. The first one maps $\textbf{X}_{n}$ to the view-independent density $\sigma_{n}$ and a corresponding feature vector. The second MLP takes the feature vector and d to produce view-dependent color $\textbf{c}_{n}$. The final color $C(\textbf{r})$ of an image pixel are rendered using volume rendering equations:
$C(\textbf{r})$ $=\int_{t_{\text{near}}}^{t_{\text{far}}}\mathcal{T}(t_{n})\sigma_{n}\textbf{c}_{n}dt_{n},$ (2)
where $\mathcal{T}(t_{n})$ is the transmittance expressed as $\exp(-\int_{t_{\text{near}}}^{t_{n}}\sigma_{k}dk)$.
@@ 3.3 Grid Sampler
[…]
@@ 3.4 Neural Acoustic Field
The goal of our neural acoustic field (NAcF), presented in Figure 4, is to learn a continuous neural representation of the scene’s acoustics. To this end, NAcF maps microphone coordinates, $\textbf{X}_{m}=(x_{m},y_{m},z_{m})$, and orientations, $\textbf{d}_{m}=(\theta_{m},\phi_{m})$, along with source position, $\textbf{X}_{s}=(x_{s},y_{s},z_{s})$, and orientation, $\textbf{d}_{s}=(\theta_{s},\phi_{s})$, to the short-time Fourier transform (STFT) of a room impulse response:
@@
We learn RIRs through their STFT representation, resulting in $\text{STFT}\in\mathbb{R}^{C\times F\times T}$, where C, F and T denote the number of channels, frequency bins, and time bins, respectively. This is motivated by the fact that the time-frequency representation is smoother and more suitable to neural network prediction than the time domain (Luo et al., 2022). Thus, NAcF learns to infer log-magnitude STFTs. To better capture perceptual effects such as reverberation, we query NAcF with STFT time bins $t$. Consequently, NAcF outputs the $F$ frequency bins corresponding to the time query. We normalize $t$ between $[0,1]$ and encode it with positional encoding. Similar to NeRF (Mildenhall et al., 2020) pose, microphone and source coordinates as well as direction are also embedded to a higher dimensional space using respectively positional encoding and spherical harmonic encoding.
@@ 3.5 Learning Objective
[…]
@@
where $\hat{s}$ is the predicted STFT, $s$ the ground-truth STFT, $\lVert.\rVert_{F}$ denotes the Frobenius norm, $\lVert.\rVert_{2}$ the L2 norm and $\epsilon=10^{-3}$. The values of our loss hyperparameters are found empirically and given in Appendix C. The spectral convergence loss emphasizes on spectral peaks helping especially in early phases of training while the spectral loss focus on small amplitude components which tends to be more important towards the later phases of training. Our loss differs from (Yamamoto et al., 2019; 2020) as we found out that the MSE spectral loss leads to a better trade-off between metrics than the L1 spectral loss (Su et al., 2022) and the MSE loss on magnitude STFT (Luo et al., 2022; Liang et al., 2023a). NeRAF results with those three losses are presented in Section 4.3. The complete learning objective can be resumed as:
@@ 4 Experiments
3
First, we benchmark NeRAF against related works (Luo et al., 2022; Su et al., 2022; Liang et al., 2023a; b) on the SoundSpaces and RAF datasets. Next, we demonstrate that cross-modal learning of the acoustic and radiance fields is also beneficial to vision, resulting in enhanced results for large complex scenes trained with sparse observations. Furthermore, we evaluate NeRAF’s ability to generalize across multiple scenes. To assess the effectiveness of our method when trained with a limited number of audio recordings, we conduct a few-shot experiment. Finally, we investigate the impact of the 3D grid and our loss combination on NeRAF’s performance. Video demonstrations showcasing NeRAF’s audio-visual generation in both real and synthetic environments are available on our project page.
4.1 Datasets
SoundSpaces.
RAF.
RWAVS.
@@ 4.2 Evaluation
Following (Su et al., 2022; Liang et al., 2023a), we assess the quality of the predicted impulse responses using Reverberation Time (T60), Clarity (C50) and Early Decay Time (EDT). T60 reflects the overall sound decay within a room by measuring the time an impulse response takes to decay by $60\,\text{dB}$. C50 measures the energy ratio between the first $50\,\text{ms}$ of RIR and the remaining portion. It relates to speech intelligibility, and the clarity of acoustics. EDT focuses on early reflections. For RAF dataset, we also evaluate the STFT error which is the absolute error between the predicted and the ground-truth log-magnitude STFTs (Défossez et al., 2018; Luo et al., 2022; Chen et al., 2024). To assess the effect of cross-modal learning on novel view synthesis, we employ commonly used metrics: PSNR, SSIM, and LPIPS.
@@
For RAF, we follow previous works experimental setup: we keep $80\%$ of data for training and $20\%$ for evaluation. Audios are cut to $0.32$ s, sampled to $48$ kHz and processed into STFT using $N_{\text{FFT}}=1024$, $W=512$ and $H=256$. For the vision part, we randomly select one third of the data from cameras 20 to 25. Nerfstudio automatically keep 90% of them for training and 10% for evaluation.
To ensure statistically significant results, we conduct each experiment 6 times and average results. More details on the implementation and architecture are presented in Appendix B and Appendix C.
@@ Results.
As SoundSpaces is a simulated dataset that presents limitations in term of visual and acoustics realism, we also present results on RAF in Table 1 (Right). NeRAF outperforms previous works best results, achieving -$\text{-}6.7$% T60, -$\text{-}14.2$% C50, -$\text{-}4.5$% EDT and -$\text{-}52.5$% STFT error. The second part of the table contains a comparison of methods using an additional energy decay loss term (Majumder et al., 2022; Liang et al., 2023b). While NACF initially relies on this term, other methods, with this new term, are denoted by ”++”. NeRAF++ still achieves state-of-the-art results, in particular for C50, EDT and STFT error. We present details about this loss in Appendix C. We showcase qualitative results of audio synthesis in Figure 5 and in Appendix I.
@@
Previous works evaluated on SoundSpaces rely on a significant amount of RIRs, which is often unavailable in real-life scenarios. Therefore, we assess the performance of our model when trained with fewer data. We conduct a few-shot experiment on one small, one medium, and one large room and average the results. As shown in Figure 7, NeRAF is better than AV-NeRF on all metrics when trained with 75% of the data. It still outperforms AV-NeRF on EDT when trained with only 25% of the data. On RAF dataset – which is a real-world dataset that already contains a more realistic amount of training data – NeRAF still outperforms the previous methods when trained with 25% less data. Results on RAF are presented in Appendix E.
@@ Conclusion.
Acknowledgments
This work was supported by the French Agence Nationale de la Recherche (ANR), under grant ANR22-CE94-0003 and was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015475 made by GENCI. We would like to thank Simon de Moreau, Raoul de Charette and the anonymous reviewers for their insightful comments and suggestions.
@@ Appendix B Architectures
The grid sampler creates a 3D grid of voxels with a given resolution. Grid coordinates are comprised between $0,1]^{3}$ fitting the scene contraction performed by NeRF. Consequently, knowing the exact dimensions of the room is not necessary. Voxels center coordinates and 18 viewing directions are queried to NeRF with no architecture modification. The grid is filled with color and density information along with voxels coordinates.
@@
NAcF consists of 2 MLP blocks. The first MLP block takes as input the multimodal fusion of grid features, encoded positions, and directions obtained through vector concatenation. It comprises 5 layers, each followed by a Leaky ReLU activation with a slope of 0.1, and outputs a 512 intermediate representation of the acoustic field. The second MLP block takes this intermediate representation as input and predicts the $F$ frequencies corresponding to the time query, aiming to learn the microphone specifics (e.g., HRTF for binaural microphones in SoundSpaces). It consists of one separate MLP head per microphone channel. The final activation layer is a tanh function scaled between [10,10][-10,10] to fit the STFT log-magnitude range.
@@ Appendix C Implementation Details
$\mathcal{L}_{A}=\lambda_{SC}\mathcal{L}_{SC}+\lambda_{SL}\mathcal{L}_{SL}+\lambda_{ED}\mathcal{L}_{ED}.\\$ (10)
@@
We follow Liang et al. (2023b) formula. Given a magnitude STFT, $\mathbf{M}$, we first compute its energy in each time window by calculating the magnitude’s square and aggregating it along the frequency dimension:
$\mathbf{M}^{\prime(d)}=\sum_{f=1}^{F}\left(\mathbf{M}^{(f,d)}\right)^{2},\ 1\leq d\leq D\kern 5.0pt,$ (11)
$\mathbf{M}^{\prime\prime(d)}=\sum_{i=d}^{D}\mathbf{M}^{\prime(i)},\ 1\leq d\leq D\kern 5.0pt,$ (12)
@@
$\text{T60}(w,w_{\text{GT}})=\frac{|\text{T60}(w)-\text{T60}(w_{\text{GT}})|}{\text{T60}(w_{\text{\text{GT}}})},$ (14)
$\text{C50}(w,w_{\text{GT}})=|\text{C50}(w)-\text{C50}(w_{\text{GT}})|,$ (15)
@@
$\text{STFT error }(\mathbf{M},\mathbf{M}_{GT})=|\log(\mathbf{M})-\log(\mathbf{M}_{GT})|$ (17)
@@ Appendix D Scene Views
Real Acoustic Fields dataset captures high-quality and dense room impulse response data paired with multi-view images using VR-NeRF ”Eyeful Tower” camera rig (Xu et al., 2023). It contains 22 cameras at different heights and pointing at different directions. We chose cameras 20 to 25 for training NeRAF’s vision part. This choice ensures that we have sufficient overlapping views. The images have a 0.7 megapixels with resolution (i.e., $684\times 1024$). For more details on those data we invite readers to refer to VR-NeRF paper. To train NeRAF, as we want to limit the number of images necessaries, we randomly sub-select one third of those images. We let Nerfstudio select 90% of them for training and 10% for evaluation.
@@ Appendix I Additional Results Visualization
Previous works (Liang et al., 2023a; b) rely on images (2D) to provide scene information to the acoustic field but we argue that conditioning on 3D information better suits the properties of sound propagation. For example, features like reverberation (linked to the room volume and surfaces) depend on complete 3D volume, not just what is visible in front of the microphone or camera. By leveraging 3D information, acoustic fields can capture the acoustics of the scene more effectively.
In this paper, we propose to use NeRF as a convenient way for providing 3D scene priors – geometry and appearance – directly to the acoustic field without needing explicit 3D annotations, such as meshes or point clouds, which are complex to obtain.
@@
As shown in Table 2 and Figure A6, joint training improves image generation in large, complex scenes with sparse visual observations. Here, we asses the impact of separate training on the audio modality. In this scenario, a NeRF (without NAcF) is fully trained and used to populate a 3D grid. Subsequently, this grid is used to train NAcF without the NeRF component. Audio generation performance remains comparable for both approaches, which is expected since the acoustic field in both setups leverages a 3D grid representation of the scene. Regarding image generation performance, for smaller scenes with sufficient observations, such as office 4, there is no notable difference in visual performance between joint and separate training. Detailed results can be found in Table A5.
@@
To supplement the results presented in Table 3, we provide an evaluation of the multi-scene model on each individual scene. The model was trained across three scenes: office 4, frl 4, and apartment 2, and evaluated on each of these scenes. We compare these results with those obtained from single-scene training in Table A6. Each grid is encoded into 1,024 features using ResNet3D.
Table A6: Comparison of NeRAF performances when trained on multiple scenes against single scene.